OpenStudy (anonymous):

find the zeros f(x)=3x^3+2x^2-8x

OpenStudy (jkristia):

factor out an x, and you have a zero at 0, and a quadratic equation x(....) = 0

OpenStudy (anonymous):

3x^3+2x^2-8x x(3x^2 + 2x - 8)=0 x((x+2)(3x-4))=0 x=0 x=-2 x=4/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw